OpenStudy (shaleiah):

A quadratic function is given as f (x) = 3x^2 + 24x + 25. Write the vertex form of the given function.

OpenStudy (shaleiah):

@563blackghost

OpenStudy (shaleiah):

@jagr2713

OpenStudy (shaleiah):

@Kainui

OpenStudy (shaleiah):

@ParthKohli

Parth (parthkohli):

Do you know how to complete the square?

OpenStudy (shaleiah):

Yes

OpenStudy (shaleiah):

x=-b/2a

Parth (parthkohli):

Oh yeah, that sounds good to me. You want something in the form\[f(x) = \left(x - (-b/2a)\right)^2 + \text{some constant here}\]

OpenStudy (shaleiah):

So, 3(x+4)^2....how do I find the rest?

OpenStudy (shaleiah):

@MarrBear

OpenStudy (anonymous):

So, Vertex would be (-b/2a, k) You would find K after you find -b/2a. So for example, you would do -24/2(3) to find your first number, then plug that into x to find your second one.

OpenStudy (shaleiah):

I got 169 :(

OpenStudy (anonymous):

Okay, lemme see if i can break it down a little more.

OpenStudy (anonymous):

Find your first point in the (h,k) Use -b/2a to find H.

OpenStudy (shaleiah):

-36

OpenStudy (anonymous):

Nope, try again. It would be -24/3(2)

OpenStudy (anonymous):

3 is your A. 24 is your B. 25 is your C.

OpenStudy (shaleiah):

h=4

OpenStudy (shaleiah):

@MarrBear

OpenStudy (anonymous):

Okay good. So you have (4, k). Input 4 in X to find your H.

OpenStudy (shaleiah):

-23

OpenStudy (anonymous):

Tell me how you got that.

OpenStudy (shaleiah):

3(-4)^2+24(-4)+25=-23 3(4)^2+24(4)+25=23

OpenStudy (shaleiah):

oops 169

OpenStudy (anonymous):

Good! So you have your points (-4,-23)

OpenStudy (shaleiah):

Thank you for your help!

OpenStudy (anonymous):

No problem!